An investment costs $3 today. It will pay two dividends of $0.31, respectively in 6 months and in 1 year. If the prevailing interest rate is 5.52% per annum, what must the final price be at the end of 1 year to make this a fair investment with an NPV of $0 ?

What should a 6-year $1,000 7.6% semi-annual coupon bond sell for, if the prevailing interest rate is 3.5%?
Answer #1

Let the final price at year end 1 be x

0 = 0.31*PVF(2.76%, 1 period) + 0.31*PVF(2.76%, 2 periods) +x* PVF(2.76%, 2 periods) -3

0 = 0.31*0.973 + 0.31*0.947 + x*0.947-3

X = $2.54

Selling price of bond should be = 38*PVAF(1.75%, 12 periods) + 1,000*PVF(1.75%, 12 periods)

= 38*10.740 +1,000*0.812

= $1,220.12